Write the net ionic equation for the precipitation reaction that occurs when aqueous solutions of potassium sulfide and aluminum sulfate are combined.


Be sure to specify states such as (aq) or (s). 


1
Expert's answer
2022-02-15T10:41:02-0500

A white gelatinous precipitate of aluminium hydroxide, Al(OH)3, is formed.


Al(3+) (aq) + 3OH(-) (aq) = Al(OH)3 (s)


Reason;


Sodium sulphide being a salt of a strong base (NaOH) and a weak acid (H2S) undergoes anionic hydrolysis in water to produce hydroxide ions as per the following equilibrium.


S(2-) (aq) + H2O (l) = HS(-) (aq) + OH(-) (aq)


Now, the aluminium ions combine with the hydroxide ions to form insoluble aluminium hydroxide.